On July 1, 2017, Lilly paid $90,000 for 450 shares of Jennifer Corp common stock. Lilly...

  1. On July 1, 2017, Lilly paid $90,000 for 450 shares of Jennifer Corp common stock. Lilly received a nontaxable stock dividend of 50 new common shares on December 15, 2019. On December 20, 2019, Lilly sold the 50 new shares for $11,000. How much should Lilly report in her 2019 return as short-term or long-term capital gain?
  1. $0
  2. $1,000
  3. $2,000
  4. $11,000
  1. McDonald, a cash basis taxpayer, is the owner of a house with two identical apartments. He resides in one apartment and rents the other apartment to a tenant under a five-year lease dated March 1, 2016, and expiring on February 28, 2021. The tenant made timely monthly rental payments of $500 for the months of January through November 2018. Rents for December 2018 and January 2019 were paid by the tenant on January 5, 2019. The following additional information for 2018 was available.

Fuel and utilities $3,600

Depreciation of building $3000

Maintenance and repairs (rental apartment) $400

Insurance on building $1,600

What amount should McDonald report as net rental income for 2018?

  1. $2,200
  2. $2,000
  3. $1,700
  4. $1,000
Answer #1

Purchase of 450 shares at $90000 in 2017.

New shares purchased has been sold in 2019

Hence when shares are sold , then it will result in a short term capital gain

Price per share = $90000/450 = $200

Purchase price of 50 shares = $200 * $50 = $10000

Sale price of 50 shares in 2019 = $11000

Hence capital gain is $1000. Correct option is b. $1000

Part 2:

Net income calulation:

Rent for 2018 = 500 * 12 = 6000

Less expenses:

Depreciation 3000

Maintanence 400

Insurance 1600

Total net income $1000